# Disentangled Dynamic Intrusion Detection

### v2

@@ II-C Dynamic Graph Convolution Networks
Dynamic graph convolution networks (GCNs) focus on evolving graph streams. There is a line of early studies in GCNs on dynamic graphs, which incorporate temporal information into graphs. These methods can be categorized into the spatio-temporal decoupled ones [53, 54, 39] and the spatio-temporal coupled ones [40, 55, 56, 57]. The former employs two separate modules to capture temporal and spatial information, e.g., DynGCN[58] performs spatial and temporal convolutions interleaved, updating model parameters to adapt to new graph snapshots, DIDA[59] Handles spatio-temporal distribution shifts by discovering invariant patterns and using intervention mechanisms to eliminate spurious impacts. The latter incorporates spatial-temporal dependencies by proposing a synchronous modeling mechanism, e.g., AST-GCN[60] adapts the graph structure and convolutional filters based on the temporal context to capture the evolving relationships in dynamic graphs, TD-GCN[61] uses temporal difference learning to update spatial-temporal graph convolutional filters, capturing the changes in graph structure and node states over time. Our DIDS is mainly inspired by the GIND [56], which adaptively aggregates information via a non-linear diffusion method. The key difference between our GCN approach and GIND is: we introduce the non-linear graph diffusion method into a multi-layer graph that considers the network topology for dynamic intrusion detection.
II-D Few-shot Learning
Few-shot learning is a machine learning paradigm designed to enable a model to learn from a small number of training instances and generalize effectively to unseen data. The existing few-shot learning methods can be divided into two categories: meta-learning-based ones [62, 63, 64], including meta-learning-based NIDS approaches [65, 66, 67, 68], and similarity-based ones. Meta-learning-based methods focus on learning a model that can adapt to new tasks with limited data. These methods often involve a meta-training phase where the model learns to learn from a variety of tasks, followed by a meta-testing phase where the model rapidly tunes to a new task. In contrast, the similarity-based ones rely on measuring the similarity or distance between the query instances and the support set to make predictions without necessitating vast types of few-shot samples. The similarity-based ones can also be grouped into augmentation-based ones [69, 70, 71, 72, 73], metric learning-based ones [74, 75, 76, 77, 32], and other similarity-based few-shot learning methods [78, 79, 80, 81]. Recent popular few-shot learning methods mainly focused on utilizing multiple learning metrics, or prototype completion. The most related to us is BSNet [27]. Ours differs from it in two aspects: first, we emphasize the importance of preserving multi-scale intrinsic representation information, rather than optimizing the original-scale information; second, we propose a joint optimization scheme to generate separated and disentangled representations to make few-shot threats distinguishable.
Finally, this work is an extension version form [20]. In [20], we address the performance inconsistency issue caused by entangled features on sufficient training data. More practical scenarios is collecting attack traffic with limited instances. It is more challenging compared to [20] with involving double optimization goals. To this end, this paper proposes an alternating optimization framework MFL to address the entangled representations in few-shot traffic threats with rigorous derivation. The proposed MFL is also a plug-and-play few-shot intrusion detection module that is compatible with other NIDS methods.
III PRELIMINARIES
III-A Multi-Layer Graphs
@@
To model the sophisticated traffic network topology, we formally define the multi-layer graphs. First, we consider a single-layer network modeled by a graph $\mathrm{G}=(\mathrm{V},\mathrm{E},\omega)$, where $V$ is the set of nodes and $E\subset\mathrm{V}\times\mathrm{V}$ is the set of edges. Here $\omega:\mathrm{V}\times\mathrm{V}\mapsto\mathrm{R}$ is an edge weight function such that each edge $e_{uv}\in\mathrm{E}$ has a weight $\omega_{uv}$. Then the multi-layer graph can be defined as follows:
@@
$\mathbb{A}=\left(\begin{array}[]{ccccc}A_{(1,1)}&\cdots&A_{(1,k)}&\cdots&A_{(1,m)}\\ \vdots&\ddots&\vdots&\ddots&\vdots\\ A_{(l,1)}&\cdots&A_{(k,k)}&\cdots&A_{(l,m)}\\ \vdots&\ddots&\vdots&\ddots&\vdots\\ A_{(m,1)}&\cdots&A_{(m,k)}&\cdots&A_{(m,m)}\end{array}\right).$ (1)
@@
III-B Edge Construction
We have constructed a multi-layer dynamic graph by defining the devices as nodes and the communications between two devices as edges. To construct the edge relationships in the multi-layer graph, we first transform the Netflow data to edges by the following definition:
$\mathbf{E}_{ij}(t)=(v_{i},l_{i},v_{j},l_{j},t,\Delta t,\mathbf{F}_{ij}(t)).$ (2)
@@ III-B Edge Construction
III-C Problem Formulation
III-C1 Dynamic Traffic Intrusion Detection
@@
III-C2 Few-shot Intrusion Detection
Given a few-shot intrusion detection task $T=\{C_{\text{tr}},C_{\text{te}}\}$, where $C_{\text{tr}}$ is a training set with few-shot traffic, and $C_{\text{te}}$ is a test set. Our goal is to design a few-shot module to accurately classify the samples in $C_{\text{te}}$ by few-shot learning on $C_{\text{tr}}$. $C_{\text{tr}}$ includes two key components, support set and query set. The former provides the few-shot traffic information to the model, and the latter contains new instances for model evaluation and optimization. The query set can be sampled from the support set for the traffic-limited scenarios. Specifically, we conduct $\mathcal{N}$-way $\mathcal{K}$-shot intrusion detection, $\mathcal{N}$-class traffic with $\mathcal{K}$ samples, i.e., $5$, in the support set. The few-shot intrusion detection model aims to conduct accurate detection in the query set, while the only available reference is the few-shot traffic in the support set.
IV MODEL
In this section, we present the DIDS-MFL, which consists of five main modules. Figure 5 shows the architecture of our proposed model. Next, we dive into the details of these modules.
Figure 5: Overview of the proposed DIDS-MFL, which consists of five modules. 1) Edge construction module builds edges based on traffic flow. 2) Statistical disentanglement module differentiates values in vectors to facilitate the identification of various attacks. 3) Representational disentanglement module learns to highlight attack-specific features. 4) Multi-Layer graph diffusion module fuses the network topology for better aggregation over evolving dynamic traffic. 5) Multi-scale few-shot learning module aims to few-shot traffic threats detection. Finally, traffic classifier takes the traffic representation as an input to yield the detection results. DIDS takes the first four steps and flows through the yellow arrows to the classifier, while DIDS-MFL takes the five steps by flowing through the blue arrows to the classifier.
IV-A Statistical Disentanglement
As we discussed in Section I, statistical distributions of traffic features are one of the main underlying causes of performance variations. i.e., separated distributions benefit the unknown attack identification, while entangled ones are indistinguishable and thus unable to help the NIDS to make accurate decisions. Therefore, our aim is there to disentangle the traffic features and make them distinguishable.
To separate the features of traffic without any prior knowledge, we formulate the differentiation as a constrained non-parametric optimization problem and approximate the optimal results by solving the Satisfiability Modulo Theory (SMT) [82]. We perform a min-max normalization on the edge feature $\mathbf{F}_{ij}(t)$. For convenience, we denote the normalized edge feature as $\mathcal{F}$, and $\mathcal{F}_{i}$ is the $i$-th normalized element.
@@ IV-A Statistical Disentanglement
$W_{\min}\leq\mathbf{w}_{i}\leq W_{\max}\quad(1\leq i\leq N),\quad\sum_{i=1}^{N}\mathbf{w}_{i}\mathcal{F}_{i}\leq B,$ (3)
@@
$\mathbf{w}_{i}\mathcal{F}_{i}\leq\mathbf{w}_{i+1}\mathcal{F}_{i+1}\quad(1\leq{i}\leq{N-1}).$ (4)
@@
$\begin{split}\widetilde{\mathbf{w}}=\arg\max(\mathbf{w}_{N}\mathcal{F}_{N}-\mathbf{w}_{1}\mathcal{F}_{1}-\\ \sum_{i=2}^{N-1}\left|2\mathbf{w}_{i}\mathcal{F}_{i}-\mathbf{w}_{i+1}\mathcal{F}_{i+1}-\mathbf{w}_{i-1}\mathcal{F}_{i-1}\right|).\end{split}$ (5)
@@
$\begin{split}\widetilde{\mathbf{w}}=\arg\max(\mathbf{w}_{N}\mathcal{F}_{N}-\mathbf{w}_{1}\mathcal{F}_{1}+\\ \sum_{i=2}^{N-1}2\mathbf{w}_{i}\mathcal{F}_{i}-\mathbf{w}_{i-1}\mathcal{F}_{i-1}-\mathbf{w}_{i+1}\mathcal{F}_{i+1}),\end{split}$ (6)
@@
$\left\{\begin{array}[]{lll}\mathbf{w}_{i}&\in&{\left[W_{\min},W_{\max}\right]}\\ \sum_{i=1}^{N}\mathbf{w}_{i}\mathbf{n}_{i}&\leq&B\\ \mathbf{w}_{i}\mathcal{F}_{i}&\leq&\mathbf{w}_{i+1}\mathcal{F}_{(i+1)}\\ 2\mathbf{w}_{i}\mathcal{F}_{i}&\leq&\mathbf{w}_{i-1}\mathcal{F}_{i-1}+\mathbf{w}_{i+1}\mathcal{F}_{i+1}.\end{array}\right.$ (7)
The subjection in Eq. 7 ensures variants’ ranges and the order-preserving property of generated representation. It also ensures symbols when removing the absolute value sign. The detailed explanation of why Eq. 6 can generate the disentangled representation is available in Apeendix 11.
Specifically, we employ Eq. 6 as an optimization objective, i.e., loss function, and utilize Adam optimizer to solve it. Then we can generate the disentangled edge representation $\mathbf{h}_{i,j}$, which can be expressed as $\mathbf{h}_{i,j}=\mathbf{w}\odot\mathcal{F}$, where the symbol $\odot$ represents the Hadamard product.
Equipped with the above non-parametric optimization, we can differentiate tens and hundreds of complex features of various attacks, mitigating the entangled distribution of statistical features. Such statistically disentangled features facilitate our model to be more sensitive to various attacks.
IV-B Representational Disentanglement
So far we have constructed edges and statistically differentiated the features of traffic flows $\mathbf{h}_{i,j}=\mathbf{w}\odot\mathcal{F}$. This module generates contextualized node representations $\mathbf{X}$ from edge representations $\mathbf{h}_{i,j}$ by using the temporal information. This involves three steps:
1) Generating updating messages: For an incoming traffic flow, we will build an edge or update the corresponding edge, which may lead to a dramatic change in the node representations involved in this interaction. We can update the node representations by utilizing this change. Therefore, we name the abrupt change an updating message and denote it as $\mathbf{c}(t)$. Specifically, we generate $\mathbf{c}(t)$ by incorporating historical memory and disentangled edge representation $\mathbf{h}_{i,j}$. The messages of node $i$ and $j$ are as follows:
$\mathbf{c}_{i}(t)$ $=\operatorname{Msg}\left(\mathbf{m}_{i}\left(t^{-}\right),\mathbf{m}_{j}\left(t^{-}\right),t,\Delta t,l_{i},l_{j},\mathbf{h_{i,j}}\right),$ (8)
$\mathbf{c}_{j}(t)$ $=\operatorname{Msg}\left(\mathbf{m}_{j}\left(t^{-}\right),\mathbf{m}_{i}\left(t^{-}\right),t,\Delta t,l_{i},l_{j},\mathbf{h_{i,j}}\right),$ (9)
where $\mathbf{h}_{i,j}$ is the disentangled edge representation. $\Delta t$ is the edge duration time, $l_{i},l_{j}$ is the layer marks of edge, $\mathbf{m}_{i}(t^{-})$ is the historical memory of the two interacting nodes, where $t^{-}$ is a historical time point, compared to the existing time point $t$, $\operatorname{Msg}$ is a learnable function, and we use RNN. The initial memory $\mathbf{m}_{i}(0)$ is $\mathbf{0}$, $\forall i\in V$. Specifically, RNN aims to generate the new memory message $\mathbf{c}(t)$ combining the historical memory $\mathbf{m}(t^{-})$ and disentangled representation $\mathbf{h}_{ij}$ by learnable weight matrices and gating mechanism.
@@ IV-B Representational Disentanglement
$\mathbf{m}_{i}(t)=\operatorname{Mem}\left(\mathbf{c}_{i}(t),\mathbf{m}_{i}\left(t^{-}\right)\right),$ (10)
where $\operatorname{Mem}$ is an encoder, and here we use GRU. We similarly update the memory of node $j$ via (10). Specifically, GRU introduces Reset Gate and Update Gate to better generate the updated node memory $\mathbf{m}_{i}(t)$ by fusing historical memory $\mathbf{m}_{i}(t^{-})$ and new memory message $\mathbf{c}_{i}(t)$. GRU considers both long-term and short-term dependencies, thus mitigating the vanishing gradient issue in temporal models.
@@
3) Generating second disentangled node representations: We can generate the node representation by utilizing the updated memory in (10), and the representation of node $i$ in time $t$ can be expressed as $\mathbf{x}_{i}(t)=\mathbf{x}_{i}(t^{-})+\mathbf{m}_{i}(t)$, where $\mathbf{{x}}_{i}(t^{-})$ is the historical representation of node $i$. In this way, we obtain the dynamic node representations of traffic in an evolving time flow. The representation of node $j$ can also be generated in a similar way. The initial value of node representations $\mathbf{X}(0)$ is $\mathbf{0}$.
@@
We aim to preserve the disentangled property in node representations. However, the update operations above may entangle them at the element level again. Therefore, we propose the second representational disentanglement, which aims to highlight the attack-specific features in node representations in the following end-to-end manner. It also ensures that the element-wised representations are close to orthogonal.
$\mathcal{L}_{\text{Dis }}=\frac{1}{2}\left\|\mathbf{X}(t){\mathbf{X}({t^{-}})}^{\top}-\mathbf{I}\right\|_{F}^{2}.$ (11)
@@
The above regularization encourages the model to learn smaller coefficients between every two elements of node representations. Such representations can be more differentiated. As the module is supervised by the signal of a specific attack, it can learn to highlight the attack-specific features, so as to improve the detection accuracy, as discussed in Section I. By doing so, we are able to mitigate the entangled distribution of representational features as mentioned at the beginning.
IV-C Multi-Layer Graph Diffusion
@@
So far, we have generated the dynamic disentangled node representations with temporal information. For further fusing the multi-layer topological structure information, we propose a multi-layer graph diffusion module. Please note that we still preserve the disentangled property by the following customized designs.
@@
We utilize the following graph diffusion method to fuse the topological information in evolving graph streams, which can capture the fine-grained spatial-temporal coupled information. The previous dynamic intrusion detection methods may lose this information in separated time gaps, due to the employed time-window or snapshots-based methods. More discussions on the ‘dynamic’ and superiority of the proposed graph diffusion approach compared to the previous methods, are available in Appendix 14.2.
$\left\{\begin{array}[]{l}\partial_{t}\mathbf{X}=\mathbf{F}\left(\mathbf{X},\mathbf{\Theta}\right)\\ \mathbf{X}(0)=\mathbf{0},\end{array}\right.$ (12)
@@
where $\mathbf{F}$ is a matrix–valued nonlinear function conditioned on graph $\mathcal{G}$, and $\mathbf{\Theta}$ is the tensor of trainable parameters. The above Eq. 12 establishes the foundation for spatial-temporal coupled information modeling.
@@
Specifically, we aim to amplify the important dimensions of disentangled representations and depress the influence of trivial feature elements in the diffusion process. To formulate this process, we consider PM (PeronaMalik) diffusion [83], a type of nonlinear filtering. It can be expressed as:
@@
$\left\{\begin{array}[]{ll}\frac{\partial x(u,t)}{\partial t}&=\operatorname{div}[g(|\nabla x(u,t)|)\nabla x(u,t)]\\ x(u,0)&=0,\end{array}\right.$ (13)
@@
Before formally formulating the multi-layer graph diffusion, we propose the following spatial-temporal influence coefficient $s_{ij}\in\mathbf{S}$ between nodes i,j in $t_{ij}$ time. The coefficient matrix considers the information changes over spatial-temporal coupled traffic data. Specifically, different layers and topology structures over traffic networks and the traffic feature interactions at different times will influence the node representations dynamically:
$s_{ij}$ $=f(l_{i}||l_{j}||\phi(t-t_{ij})),$ (14)
$f(\mathbf{x})$ $=\mathbf{W}^{(2)}\cdot\operatorname{ReLU}\left(\mathbf{W}^{(1)}\mathbf{x}\right),$ (15)
where ||||| | is a concatenate operator, $\phi(\cdot)$ is a generic time encoder [84] to generate temporal representations, $\mathbf{W}^{(1)}$ and $\mathbf{W}^{(2)}$ are the parameters of the first and second layer MLP.
Now we formally propose the multi-layer graph diffusion module. We first define a differential operator on the multi-layer graph, aiming to transfer the above continuous PM diffusion in Eq. 13 to multi-layer graphs. As known from previous literature [85], the gradient operator corresponds to the instance matrix $\mathbf{M}$, while the divergence operator corresponds to the matrix $\mathbf{M}^{\top}$, and we can compute the matrix $\mathbf{M}$ by the equation $\mathbf{M}^{T}\mathbf{M}=\mathbf{D}-\mathbf{A}$, where $\mathbf{D}$ is the diagonal matrix. Then our novel multi-layer diffusion can be expressed as:
${\partial\mathbf{X}_{t}}=-\mathbf{M}^{\top}\sigma(\mathbf{M}\mathbf{X}\mathbf{K}^{\top})\mathbf{S}\left(\mathbf{M}\mathbf{X}\mathbf{K}^{\top}\right)\mathbf{K},$
where $\mathbf{K}$ is a transformation matrix, $\mathbf{S}$ is the structure-temporal influence coefficients calculated in Eq. 14 and $\sigma(\cdot)$ represents the function $\exp(-|\cdot|)$. The solution to equation 16 can be expressed as:
$\mathbf{X}_{t+\Delta t}=\mathbf{X}_{t}+\int_{t}^{t+\Delta t}{\partial_{t}\mathbf{X}_{t}}d\tau,$ (17)
@@
IV-D Classifier and Loss Function
For DIDS, we make the two-step predictions for intrusion detection. We utilize the first MLP to classify whether the traffic is benign or anomalous and utilize the second MLP to detect the specific type of attack. When an unknown attack invades, the direct multi-classifications will be easy to fail to assign the anomalous label thus leading to poor performance. In contrast, through the first-step binary classification, DIDS will focus more on inconsistencies with normal behavior to improve the performance of detecting unknown attacks. The following second multi-classification will further alert the administrators that what kind of attack it is more similar to so that similar mitigation measures can be taken. Specifically, the intrusion loss can be expressed as:
$\mathcal{L}_{\text{Int}}=-\sum_{i=1}^{m}(\log(1-p_{\mathrm{nor},i})+\log(p_{\mathrm{att},i})+\sum_{j=1}^{K}y_{i,k}\log(p_{i,k}))$

### v3

@@ 2.3. Dynamic Graph Convolution Networks
Dynamic graph convolution networks (GCNs) focus on evolving graph streams. There is a line of early studies in GCNs on dynamic graphs, which incorporate temporal information into graphs. These methods can be categorized into the spatio-temporal decoupled ones (Pareja et al., 2020; Kalander et al., 2020b; Cong et al., 2021) and the spatio-temporal coupled ones (Song et al., 2020; Poli et al., 2019; Gu et al., 2020; Chen et al., 2022; Chamberlain et al., 2021). The former employs two separate modules to capture temporal and spatial information respectively, and the latter ones incorporate spatial-temporal dependencies by proposing a synchronous modeling mechanism. Our 3D-IDS is mainly inspired by the GIND (Chen et al., 2022), which adaptively aggregates information via a non-linear diffusion method. The key difference between our GCN approach and GIND is: we introduce the non-linear graph diffusion method into a multi-layer graph that considers the network topology for dynamic intrusion detection.
3. PRELIMINARIES
3.1. Multi-Layer Graphs
@@
We consider a single-layer network modeled by a graph $\mathrm{G}=(\mathrm{V},\mathrm{E},\omega)$, where $V$ is the set of nodes and $E\subset\mathrm{V}\times\mathrm{V}$ is the set of edges. Here $\omega:\mathrm{V}\times\mathrm{V}\mapsto\mathrm{R}$ is an edge weight function such that each edge $e_{uv}\in\mathrm{E}$ has a weight $\omega_{uv}$. A multi-layer network can be modeled by a multi-layer graph:
@@
(1) $\mathbb{A}=\left(\begin{array}[]{ccccc}A_{(1,1)}&\cdots&A_{(1,k)}&\cdots&A_{(1,m)}\\ \vdots&\ddots&\vdots&\ddots&\vdots\\ A_{(l,1)}&\cdots&A_{(k,k)}&\cdots&A_{(l,m)}\\ \vdots&\ddots&\vdots&\ddots&\vdots\\ A_{(m,1)}&\cdots&A_{(m,k)}&\cdots&A_{(m,m)}\end{array}\right).$
@@
3.2. Edge Construction
We have constructed a multi-layer dynamic graph by defining the devices as nodes and the communications between two devices as edges. Specifically, we first transform the Netflow data to edges by the following definition:
(2) $\mathbf{E}_{ij}(t)=(v_{i},l_{i},v_{j},l_{j},t,\Delta t,\mathbf{F}_{ij}(t)).$
@@ 3.2. Edge Construction
3.3. Problem Formulation
@@
4. MODEL
In this section, we present the 3D-IDS, which consists of four main modules. Figure 3 shows the architecture of our proposed model. Next, we dive into the details of these modules.
Figure 3. Overview of the proposed 3D-IDS, which consists of five modules. 1) Edge construction module builds edges based on traffic flow. 2) Statistical disentanglement module differentiates values in vectors to facilitate the identification of various attacks. 3) Representational disentanglement module learns to highlight attack-specific features. 4) Multi-Layer graph diffusion module fuses the network topology for better aggregation over evolving dynamic traffic. 5) Traffic classifier takes the traffic representation as an input to yield the detection results.
4.1. Statistical Disentanglement
As we discussed in Section 1, statistical distributions of traffic features are one of the main underlying causes of performance variations. i.e., separated distributions benefit the unknown attack identification, while entangled ones are indistinguishable and thus unable to help the NIDS to make accurate decisions. Therefore, our aim is there to disentangle the traffic features and make them distinguishable.
To separate the features of traffic without any prior knowledge, we formulate the differentiation as a constrained non-parametric optimization problem and approximate the optimal results by solving the Satisfiability Modulo Theory (SMT) (De Moura and Bjørner, 2008). We perform a min-max normalization on the edge feature $\mathbf{F}_{ij}(t)$. For convenience, we denote the normalized edge feature as $\mathcal{F}$, and $\mathcal{F}_{i}$ is the $i$-th normalized element.
@@ 4.1. Statistical Disentanglement
(3) $W_{\min}\leq\mathbf{w}_{i}\leq W_{\max}\quad(1\leq i\leq N),\quad\sum_{i=1}^{N}\mathbf{w}_{i}\mathcal{F}_{i}\leq B,$
@@
(4) $\mathbf{w}_{i}\mathcal{F}_{i}\leq\mathbf{w}_{i+1}\mathcal{F}_{i+1}\quad(1\leq{i}\leq{N-1}).$
@@
(5) $\begin{split}\widetilde{\mathbf{w}}=\arg\max(\mathbf{w}_{N}\mathcal{F}_{N}-\mathbf{w}_{1}\mathcal{F}_{1}-\\ \sum_{i=2}^{N-1}\left|2\mathbf{w}_{i}\mathcal{F}_{i}-\mathbf{w}_{i+1}\mathcal{F}_{i+1}-\mathbf{w}_{i-1}\mathcal{F}_{i-1}\right|).\end{split}$
@@
(6) $\begin{split}\widetilde{\mathbf{w}}=\arg\max(\mathbf{w}_{N}\mathcal{F}_{N}-\mathbf{w}_{1}\mathcal{F}_{1}+\\ \sum_{i=2}^{N-1}2\mathbf{w}_{i}\mathcal{F}_{i}-\mathbf{w}_{i-1}\mathcal{F}_{i-1}-\mathbf{w}_{i+1}\mathcal{F}_{i+1}),\end{split}$
@@
(7) $\left\{\begin{array}[]{lll}\mathbf{w}_{i}&\in&{\left[W_{\min},W_{\max}\right]}\\ \sum_{i=1}^{N}\mathbf{w}_{i}\mathbf{n}_{i}&\leq&B\\ \mathbf{w}_{i}\mathcal{F}_{i}&\leq&\mathbf{w}_{i+1}\mathcal{F}_{(i+1)}\\ 2\mathbf{w}_{i}\mathcal{F}_{i}&\leq&\mathbf{w}_{i-1}\mathcal{F}_{i-1}+\mathbf{w}_{i+1}\mathcal{F}_{i+1}.\end{array}\right.$
For the above formation, we can generate the disentangled edge representation $\mathbf{h}_{i,j}$ by solving the above problem, which can be expressed as $\mathbf{h}_{i,j}=\mathbf{w}\odot\mathcal{F}$, where the symbol $\odot$ represents the Hadamard product.
Equipped with the above non-parametric optimization, we can differentiate tens and hundreds of complex features of various attacks, mitigating the entangled distribution of statistical features. Such statistical disentangled features facilitate our model to be more sensitive to various attacks.
4.2. Representational Disentanglement
So far we have constructed edges and statistically differentiated the features of traffic flows. This module generates contextualized node representations $\mathbf{X}$ from edge representations. This involves three steps:
1) Generating updating messages: For an incoming traffic flow, we will build an edge or update the corresponding edge, which may lead to a dramatic change in the node representations involved in this interaction. We can update the node representations by utilizing this change. Therefore, we name the abrupt change an updating message and denote it as $c(t)$. Specifically, we generate $c(t)$ by incorporating historical memory and disentangled edge representation $\mathbf{h}_{i,j}$. The messages of node $i$ and $j$ are as follows:
(8) $\mathbf{c}_{i}(t)$ $=\operatorname{Msg}\left(\mathbf{m}_{i}\left(t^{-}\right),\mathbf{m}_{j}\left(t^{-}\right),t,\Delta t,l_{i},l_{j},\mathbf{h_{i,j}}\right),$
(9) $\mathbf{c}_{j}(t)$ $=\operatorname{Msg}\left(\mathbf{m}_{j}\left(t^{-}\right),\mathbf{m}_{i}\left(t^{-}\right),t,\Delta t,l_{i},l_{j},\mathbf{h_{i,j}}\right),$
where $\mathbf{h}_{i,j}$ is the disentangled edge representation. $\Delta t$ is the edge duration time, $l_{i},l_{j}$ is the layer marks of edge, $m_{i}(t^{-})$, $m_{j}(t^{-})$is the historical memory of the two interacting nodes, $\operatorname{Msg}$ is a learnable function, and we use RNN. The initial memory $\mathbf{m}(0)$ is set as $\mathbf{0}$.
@@ 4.2. Representational Disentanglement
(10) $\mathbf{m}_{i}(t)=\operatorname{Mem}\left(\mathbf{c}_{i}(t),\mathbf{m}_{i}\left(t^{-}\right)\right),$
where $\operatorname{Mem}$ is an encoder, and here we use GRU. We also update the memory of node $j$ via (10).
@@
3) Generating second disentangled node representations: We can generate the node representation by utilizing the updated memory in (10), and the representation of node $i$ in time $t$ can be expressed as $\mathbf{x}_{i}(t)=\mathbf{x}_{i}(t^{-})+\mathbf{m}_{i}(t)$, where $\mathbf{{x}}_{i}(t^{-})$ is the historical representation of node $i$. The representation of node $j$ can also be generated in a similar way. The initial value of node representations $\mathbf{X}(0)$ is $\mathbf{0}$.
@@
We aim to preserve the disentangled property in node representations. However, the update operations above may entangle them at the element level again. Therefore, we propose the second representational disentanglement, which aims to highlight the attack-specific features for the following end-to-end training. It also ensures that the element-wised representations are close to orthogonal.
(11) $\mathcal{L}_{\text{Dis }}=\frac{1}{2}\left\|\mathbf{X}(t){\mathbf{X}({t^{-}})}^{\top}-\mathbf{I}\right\|_{F}^{2}.$
@@
The above regularization encourages the model to learn smaller coefficients between every two elements of node representations. Such representations can be more differentiated. As the module is supervised by the signal of a specific attack, it can learn to highlight the attack-specific features, so as to improve the detection scores, as discussed in Section 1. By doing so, we are able to mitigate the entangled distribution of representational features as mentioned at the beginning.
4.3. Multi-Layer Graph Diffusion
@@
So far, we have generated the disentangled node representations with temporal information. For further fusing the multi-layer topological structure information, we propose a multi-layer graph diffusion module, and please note that we still preserve the disentangled property by customized designs.
@@
We utilize the following graph diffusion method to fuse the topological information in evolving graph streams, which can capture the fine-grained continuous spatio-temporal information. The previous dynamic intrusion detection methods may lose this information in separated time gaps, due to the employed time-window or snapshots-based methods.
(12) $\left\{\begin{array}[]{l}\partial_{t}\mathbf{X}=\mathbf{F}\left(\mathbf{X},\mathbf{\Theta}\right)\\ \mathbf{X}_{0}=\mathbf{C},\end{array}\right.$
@@
where $\mathbf{F}$ is a matrix–valued nonlinear function conditioned on graph $\mathcal{G}$, and $\mathbf{\Theta}$ is the tensor of trainable parameters.
@@
Specifically, we aim to amplify the important dimensions of disentangled representations and depress the influence of trivial feature elements in the diffusion process. To formulate this process, we consider PM (PeronaMalik) diffusion (Perona and Malik, 1990), a type of nonlinear filtering. It can be expressed as:
@@
(13) $\left\{\begin{array}[]{ll}\frac{\partial x(u,t)}{\partial t}&=\operatorname{div}[g(|\nabla x(u,t)|)\nabla x(u,t)]\\ x(u,0)&=c,\end{array}\right.$
@@
In addition, the edges in dynamic multi-layer graphs always have different timestamps and are in different layers. There are also different interactions at different layers and times. These factors will dynamically affect the information exchange results over graphs. Therefore, we aim to incorporate the above factors into the diffusion process. We propose the following layer-temporal coefficient $s_{ij}$ between nodes i,j in $t_{ij}$ time:
(14) $s_{ij}$ $=f(l_{i}||l_{j}||\phi(t-t_{ij})),$
(15) $f(\mathbf{x})$ $=\mathbf{W}^{(2)}\cdot\operatorname{ReLU}\left(\mathbf{W}^{(1)}\mathbf{x}\right),$
where $f(\cdot)$ is defined in (15), $\phi(\cdot)$ represents a generic time encoder (Xu et al., 2020), $\mathbf{W}^{(1)}$ and $\mathbf{W}^{(2)}$ are the parameters of the first layer and second layer MLP.
To transfer the above continuous PM diffusion to the multi-layer graph, we need to define the differential operators on the multi-layer graph. As known from previous literature (Chung, 1997), the gradient operator corresponds to the instance matrix $\mathbf{M}$, while the divergence operator corresponds to the matrix $\mathbf{M}^{\top}$, and we can compute the matrix $M$ by the equation $\mathbf{M}^{T}\mathbf{M}=\mathbf{D}-\mathbf{A}$, where $\mathbf{D}$ is the diagonal matrix. Then our novel multi-layer diffusion can be expressed as:
(16) ${\partial\mathbf{X}_{t}}=-\mathbf{M}^{\top}\sigma(\mathbf{M}\mathbf{X}\mathbf{K}^{\top})\mathbf{S}\left(\mathbf{M}\mathbf{X}\mathbf{K}^{\top}\right)\mathbf{K},$
where $\mathbf{K}$ is a transformation matrix, $\mathbf{S}$ is the layer-temporal influence coefficients and $\sigma(x)$ represents the function $\exp(-|x|)$. The solution to equation 16 can be expressed as:
(17) $\mathbf{X}_{t+\Delta t}=\mathbf{X}_{t}+\int_{t}^{t+\Delta t}{\partial_{t}\mathbf{X}_{t}}d\tau,$
@@
4.4. Classifier and Loss Function
Finally, we make the two-step predictions for intrusion detection. We utilize the first MLP to classify whether the traffic is benign or anomalous and utilize the second MLP to detect the specific type of attack. When an unknown attack invades, the direct multi-classifications will be easy to fail to assign the anomalous label thus leading to poor performance. In contrast, through the first-step binary classification, 3D-IDS will focus more on inconsistencies with normal behavior to improve the performance of detecting unknown attacks. The following second multi-classification will further alert the administrators that what kind of attack it is more similar to so that similar mitigation measures can be taken. Specifically, the intrusion loss can be expressed as:
(18) $\mathcal{L}_{\text{Int}}=-\sum_{i=1}^{m}(\log(1-p_{\mathrm{nor},i})+\log(p_{\mathrm{att},i})+\sum_{i=1}^{m}\sum_{j=1}^{K}y_{i,k}{\log(p_{i,k}}).$